# Experimental Multi-site Testbed for Advanced Control and Optimization of Hybrid Energy Systems

### v1

@@ IV Experimental Test Case: Solar Smoothing using Physical Battery Storage
The increasing penetration of PV generation introduces rapid power fluctuations from cloud transitions and irradiance variability, which challenges grid stability and power quality. A BESS can mitigate these fluctuations through fast response and bidirectional power capability, absorbing excess power during PV spikes and injecting power during drops. By coordinating PV generation with BESS operation, HES can smooth the power delivered to the grid. Grid operators impose ramp-rate limits to keep fluctuations within acceptable ranges, defined as:
@@
where $P_{\text{PV}}(t)$ is the PV power at time $t$, $\Delta t$ is the time interval, and $P_{\text{rated}}$ is the rated PV capacity. In this study, the maximum allowable ramp rate is set to $\pm 5\%$/min, which lies within the 2-10%/min range specified in several PV grid codes studies.
To evaluate the smoothing capability at ATL, a CHIL experiment was conducted using the experimental setup shown in Fig. 4. The experiment uses historical 5-second PV data from one inverter at the HSTC site. The PV profile is scaled to match the available battery capacity; this scaling does not affect ramp-rate behavior because ramp rates are expressed relative to the rated capacity. The OPAL-RT sends PV power and measured battery voltage to the TI controller, which runs a 30-minute moving-average smoothing algorithm, detailed in Algorithm 1. The controller issues battery current setpoints to the power supply through OPAL-RT. The battery management system (BMS) estimates the battery state of charge (SOC).
Figure 4: Experimental setup for solar smoothing CHIL validation showing two series-connected Enphase batteries with a BMS, bidirectional DC power supply, TI control card, and OPAL-RT real-time simulator.
In : $P_{\mathrm{PV}}[k]$, $V_{\mathrm{batt}}[k]$
Out : $I_{\mathrm{set}}[k]$
@@
  $P_{\mathrm{buf}}[([k-1]\bmod N)+1]\leftarrow P_{\mathrm{PV}}[k]$
  // 2) Compute smoothed power
@@
  // 4) Current setpoint
  $I_{\mathrm{set}}[k]\leftarrow\dfrac{P_{\mathrm{batt}}[k]}{V_{\mathrm{batt}}[k]}$
  // 5) Output and increment
  Output $I_{\mathrm{set}}[k]$
  $k\leftarrow k+1$ ; sleep($T_{s}$)
@@
A two-hour snapshot of the CHIL experiment, shown in Fig. 5, highlights the system smoothing performance. Fig. 5(a) shows the raw PV power and compares the smoothed results from simulation and experiment, which indicate that the results align closely. Fig. 5(b) presents the dispatched battery power, and Fig. 5(c) shows the resulting SOC, which remain within safe operating limits throughout the experiment. The raw PV profile shows a maximum ramp rate of 56%/min, exceeding the $\pm 5\%$/min grid limit, whereas the BESS effectively reduces the maximum ramp rate to 3.75%/min. The corresponding distribution of ramp rates is shown in Fig. 5(d), which illustrates the significant reduction in extreme events. These results validate the capability of the ATL facility to effectively implement and test PV smoothing strategies using BESS.
Figure 5: Two-hour snapshot of HIL experimental PV power smoothing using historical 5-second HSTC data and ATL battery setup. (a) raw and smoothed PV power, here, $P_{\text{PV}}$ denotes the raw PV power, $\hat{P}_{\text{PV}}^{e}$ and $\hat{P}_{\text{PV}}^{s}$ denote the smoothed PV power from experiment and simulation, respectively, (b) dispatched battery power, (c) battery SOC, and (d) distribution of ramp rates.

### v2

@@ IV Experimental Test Case: Solar Smoothing using Physical Battery Storage
The increasing penetration of PV generation introduces rapid power fluctuations from cloud transitions and irradiance variability, which challenges grid stability and power quality. A BESS can mitigate these fluctuations through fast response and bidirectional power capability, absorbing excess power during PV spikes and injecting power during drops. By coordinating PV generation with BESS operation, HES can smooth the power delivered to the grid. Grid operators impose ramp rate limits to keep fluctuations within acceptable ranges, defined as:
@@
where $P_{\text{PV}}(t)$ is the PV power at time $t$, $\Delta t$ is the time interval, and $P_{\text{rated}}$ is the rated PV capacity. In this study, the maximum allowable ramp rate is set to $\pm 10~\%$/min, consistent with limits specified in PV grid codes [9].
Figure 3: Experimental setup for solar smoothing HIL validation showing two series-connected Enphase batteries with a BMS, bidirectional DC power supply, TI control card, and OPAL-RT real-time simulator.
Input : $P_{\mathrm{PV}}[k]$
1exOutput : $P_{\mathrm{batt,h}}[k],\ P_{\mathrm{batt,s}}[k]$
Parameters: $N,\ T_{s}$, $w_{h}\leftarrow\frac{P_{h}^{\max}}{P_{h}^{\max}+P_{s}^{\max}}$; $w_{s}\leftarrow 1-w_{h}$
$P_{\mathrm{buf}}[i]\leftarrow 0,\ i=1{:}N$; $k\leftarrow 1$
@@
  // 3) Total battery command
 $P_{\mathrm{batt,h}}[k]\leftarrow w_{h}\cdot P_{\mathrm{batt}}[k]$
  // 4) Battery power allocation
  $P_{\mathrm{batt,s}}[k]\leftarrow w_{s}\cdot P_{\mathrm{batt}}[k]$
 Output $P_{\mathrm{batt,h}}[k],\ P_{\mathrm{batt,s}}[k]$
  // 5) Output
  $k\leftarrow k+1$, sleep($T_{s}$)
@@
To evaluate the smoothing capability at ATL, an HIL experiment was conducted using the experimental setup shown in Fig. 3. The experiment uses historical 5-second PV power data from the HSTC site. To meet the power requirements for PV smoothing, the testbed integrates physical battery hardware (1.2 kW/2.4 kWh) together with three battery equivalent circuit models connected in series and implemented as the simulated battery (7.2 kW/7.2 kWh). The OPAL-RT sends PV power to the TI controller, which runs a 30-minute moving-average smoothing algorithm, detailed in Algorithm 1. An energy management system (EMS) distributes the required battery power between the two battery systems based on their maximum power ratings. The controller issues setpoints for the battery to the power supply through OPAL-RT. The battery management system (BMS) estimates the battery SOC.
Fig. 4 presents a two-hour snapshot of the HIL experiment. The raw PV power exhibits a maximum ramp rate of 56 %/min, significantly exceeding the $\pm 10~\%$/min grid limit. Fig.4(a) demonstrates that the BESS effectively reduces this to 9.45%/min. The simulation and experimental results show close agreement, with a root mean square error (RMSE) of 0.25 kW (2.4 % of maximum power). This error is attributed to communication latency, ADC quantization, sensor noise, and sampling effects. Fig. 4(b) presents dispatched battery power for the physical and simulated batteries. Fig. 4(c) shows the SOC of both batteries. The physical battery exhibits larger deviations due to its lower capacity and maximum power; however, both batteries remain within safe operating limits. Fig. 4(d) illustrates the ramp rate distribution, showing reduced extreme events and zero grid-constraint violations. These results validate ATL’s capability to implement and test PV smoothing strategies under multi-site operation.
Figure 4: Two-hour snapshot of HIL experimental PV power smoothing using 5-second HSTC data and the ATL battery setup. (a) Raw and smoothed PV power, where $P_{\text{PV}}$ denotes the raw PV power, and $\hat{P}_{\text{PV}}^{e}$ and $\hat{P}_{\text{PV}}^{s}$ denote the experimentally measured and simulated smoothed PV power, respectively, (b) dispatched battery power of the physical and simulated batteries, (c) battery SOC, and (d) ramp-rate distributions.